OpenStudy (nuccioreggie):

Patty's Pet Company charges $0.55 per pound to ship pet supplies. Part A: Write an equation to determine the total cost, c, to ship p pounds of pet supplies. Use your equation to determine the cost of shipping 2 pounds of pet supplies. (6 points) Part B: If the company reduces the cost to ship pet supplies by 0.05 per pound, write an equation to determine the total cost, c, to ship p pounds of pet supplies with the reduced cost. (4 points)

OpenStudy (nuccioreggie):

@Michele_Laino Please EXPLAIN

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

if the company requests $0.55 per pounds, then it requests: $0.55*n for n pounds

OpenStudy (nuccioreggie):

Tha the answer for part A?

OpenStudy (michele_laino):

yes! The cost for p pounds, is: 0.55*p

OpenStudy (michele_laino):

if p=2, then we have: c=0.55*2=... please complete

OpenStudy (nuccioreggie):

Can you complete it for me so i would understand better please

OpenStudy (michele_laino):

it is simple, what is 0.55*2 ?

OpenStudy (nuccioreggie):

1.1

OpenStudy (michele_laino):

ok! that's right!

OpenStudy (michele_laino):

Part B: now the new unit price is: 0.55-0.05=...?

OpenStudy (michele_laino):

please complete that computation

OpenStudy (nuccioreggie):

Ok

OpenStudy (michele_laino):

what do you get? @nuccioreggie

OpenStudy (nuccioreggie):

0.5

OpenStudy (nuccioreggie):

So can you right it out for me ease

OpenStudy (michele_laino):

perfect! so the new cost per pound, is : c= 0.50*p

OpenStudy (nuccioreggie):

Ok so now can you right the whole part B answer out for me PLEEEEASE

OpenStudy (michele_laino):

I have already answered to part B

OpenStudy (michele_laino):

the requested formula, or function, is: c=0.50*p